User Tools

Site Tools


handover:hob010

HOB010

Hobart 26m

VSN OAN+0103, start pos 1842.7 GB

  • 1730UT experiment started OK (JS)
  • 23:54 UT drive power switched off, via DESK control (remote vdesk operator unaware of current experiment?), power came back up at 23:56 UT and telescope back on scheduled source (Warren)
  • 1040UT mk5cn connection error, Halt experiment and restart mk5ho (Note start DIM on oper@mk5ho with DIMino -m0) (JS)
  • Reestablish at 1119UT
    • Scans missed 357-1040 through 357-1119
  • 1240UT repeated mk5cn errors, reset mk5ho again (JS)
    • lost scans 357-1240 through 357-1255
  • NOTE: REcording switched to module MAT+0005 at 1255UT (after mk5ho reset) 692 GB of data recorded here

Hobart 12m

VSN HOB+0097, start pos 10765.4

  • 1730UT experiment started OK (JS)
  • 2031UT Channels 10 and 12 have negative tsys (only recording x-band though??) (JS)
  • 1040UT accidental closure of DIM server while fixing Ho,fixed with full FS reset at 1105UT
    • Scans missed 357-1040 through 357-1105
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/hob010.txt · Last modified: 2016/12/22 18:08 by Lucas Hyland